# Borel 1 type mappings and the respective equi-families

### v1

@@ 3.2. Closures in the topology of pointwise convergence
([Kel55], Theorem 14, Chapter 7) Let $X$ be a topological space, $(Y,d)$ be a metric space and $\mathcal{F}$ be a family of functions from $X$ to $Y$. If $\mathcal{F}$ is equicontinuous at $x$, then the closure $\overline{\mathcal{F}}$ of $\mathcal{F}$ relative to the topology of pointwise convergence is also equicontinuous at $x$.
@@
([Lec03], Proposition 35) Let $(X,d_{X})$ and $(Y,d_{Y})$ be metric spaces. If $\mathcal{F}$ is an equi-Baire 1 family of functions from $X$ to $Y$, then the closure $\overline{\mathcal{F}}$ of $\mathcal{F}$ relative to the topology of pointwise convergence is equi-Baire 1 too.
@@
([BHH25], Proposition 4.1) Let $(X,d_{X})$ and $(Y,d_{Y})$ be metric spaces. If $\mathcal{F}$ is an equi-Lebesgue family of functions from $X$ to $Y$, then the closure $\overline{\mathcal{F}}$ of $\mathcal{F}$ relative to the topology of pointwise convergence is equi-Lebesgue too.
@@ 3.3. Applications to functions defined on $X\times Y$
We are going to solve two problems from the earlier paper [BKS22].
@@
[BKS22, Prob. 3.7] Suppose that $\mathcal{F}\subseteq\mathbb{R}^{X\times Y}$ is a family of separately equi-continuous functions defined on the product of Polish spaces $X$ and $Y$, i.e.
@@
Let us give a positive solution in a general setting. Bouziad [Bou12, Proposition 4.3] proved the following fact. If $X,Y$ are topological spaces and $Z$ is a metric space, then every separately continuous function $f\colon X\times Y\to Z$ is weakly separated. Now, we obtain the following corollary which solves the problem.
@@
Let us state the next problem [BKS22, Prob. 3.11] in a strictly formal version as follows.
@@
Let us show a negative solution of Problem 3.15. Grande noticed that there exists a function $f\colon\mathbb{R}^{2}\to\mathbb{R}$ that has all $x$-sections and all $y$-sections of class Baire 1 but $f$ is not Lebesgue measurable (or it does not have the Baire property); see e.g. [Gra83, Remark 2]. One can use this fact to obtain the respective counterexample for Problem 3.15. However, we propose a different simple construction. Let $B\subseteq\mathbb{R}$ be a Bernstein nonmeasurable set. It is known that for any uncountable Borel set $A\subseteq\mathbb{R}$ the intersection $B\cap A$ is not a Borel set. Let $g\colon\mathbb{R}^{2}\to\mathbb{R}$ be the characteristic function of the set $B^{*}=\{(x,x)\colon x\in B\}$. Then $g$ is separately Baire 1. Suppose $A\subseteq\mathbb{R}$ is a nonmeager $G_{\delta}$ set such that $g\restriction_{A\times\mathbb{R}}$ is Baire 1. Then $A\cap B$ should be Borel, a contradiction. Similarly, the function $f\colon\mathbb{R}^{2}\to\mathbb{R}^{\mathbb{N}}$ given by $f=(g,g,\dots)$ is separately Baire 1 and $f\restriction_{A\times\mathbb{R}}$ is not jointly Baire 1 for every nonmeager $G_{\delta}$ set $A\subseteq\mathbb{R}$. This by Remark 3.16 leads to the negative solution of Problem 3.15.
Grande proved in [Gra83, Thm 6] the following interesting fact on real functions of two variables. If $f\colon\mathbb{R}^{2}\to\mathbb{R}$ is a function with all $x$-sections that form a family having (PECP) and all $y$-sections of Borel class $\alpha$, then $f$ is of Borel class $\alpha$. We have observed that this result with the same proof remains valid if $f\colon X\times Y\to Z$ where $X,Y,Z$ are metric spaces and $Y$ is separable. Below, we propose a further extension of this theorem.
@@
Ackonwledgement. We would like to thank Tomek Natkaniec for showing us Theorem 6 in Grande’s article [Gra83] which is relevant to our consideration of Problem 3.15.

### v2

@@ 3.2. Closures in the topology of pointwise convergence
([22], Theorem 14, Chapter 7) Let $X$ be a topological space, $(Y,d)$ be a metric space and $\mathcal{F}$ be a family of functions from $X$ to $Y$. If $\mathcal{F}$ is equicontinuous at $x$, then the closure $\overline{\mathcal{F}}$ of $\mathcal{F}$ relative to the topology of pointwise convergence is also equicontinuous at $x$.
@@
([26], Proposition 35) Let $(X,d_{X})$ and $(Y,d_{Y})$ be metric spaces. If $\mathcal{F}$ is an equi-Baire 1 family of functions from $X$ to $Y$, then the closure $\overline{\mathcal{F}}$ of $\mathcal{F}$ relative to the topology of pointwise convergence is equi-Baire 1 too.
@@
([4], Proposition 4.1) Let $(X,d_{X})$ and $(Y,d_{Y})$ be metric spaces. If $\mathcal{F}$ is an equi-Lebesgue family of functions from $X$ to $Y$, then the closure $\overline{\mathcal{F}}$ of $\mathcal{F}$ relative to the topology of pointwise convergence is equi-Lebesgue too.
@@ 3.3. Applications to functions defined on $X\times Y$
We are going to solve two problems from the earlier paper [5].
@@
[5, Prob. 3.7] Suppose that $\mathcal{F}\subseteq\mathbb{R}^{X\times Y}$ is a family of separately equi-continuous functions defined on the product of Polish spaces $X$ and $Y$, i.e.
@@
Let us give a positive solution in a general setting. Bouziad [6, Proposition 4.3] proved the following fact. If $X,Y$ are topological spaces and $Z$ is a metric space, then every separately continuous function $f\colon X\times Y\to Z$ is weakly separated. Now, we obtain the following corollary which solves the problem.
@@
Let us state the next problem [5, Prob. 3.11] in a strictly formal version as follows.
@@
Let us show a negative solution of Problem 3.15. Grande noticed that there exists a function $f\colon\mathbb{R}^{2}\to\mathbb{R}$ that has all $x$-sections and all $y$-sections of class Baire 1 but $f$ is not Lebesgue measurable (or it does not have the Baire property); see e.g. [15, Remark 2]. One can use this fact to obtain the respective counterexample for Problem 3.15. However, we propose a different simple construction. Let $B\subseteq\mathbb{R}$ be a Bernstein nonmeasurable set. It is known that for any uncountable Borel set $A\subseteq\mathbb{R}$ the intersection $B\cap A$ is not a Borel set. Let $g\colon\mathbb{R}^{2}\to\mathbb{R}$ be the characteristic function of the set $B^{*}=\{(x,x)\colon x\in B\}$. Then $g$ is separately Baire 1. Suppose $A\subseteq\mathbb{R}$ is a nonmeager $G_{\delta}$ set such that $g\restriction_{A\times\mathbb{R}}$ is Baire 1. Then $A\cap B$ should be Borel, a contradiction. Similarly, the function $f\colon\mathbb{R}^{2}\to\mathbb{R}^{\mathbb{N}}$ given by $f=(g,g,\dots)$ is separately Baire 1 and $f\restriction_{A\times\mathbb{R}}$ is not jointly Baire 1 for every nonmeager $G_{\delta}$ set $A\subseteq\mathbb{R}$. This by Remark 3.16 leads to the negative solution of Problem 3.15.
Grande proved in [15, Thm 6] the following interesting fact on real functions of two variables. If $f\colon\mathbb{R}^{2}\to\mathbb{R}$ is a function with all $x$-sections that form a family having (PECP) and all $y$-sections of Borel class $\alpha$, then $f$ is of Borel class $\alpha$. We have observed that this result with the same proof remains valid if $f\colon X\times Y\to Z$ where $X,Y,Z$ are metric spaces and $Y$ is separable. Below, we propose a further extension of this theorem.
@@
Ackonwledgement. We would like to thank Tomek Natkaniec for showing us Theorem 6 in Grande’s article [15] which is relevant to our consideration of Problem 3.15.